a. Can [ i ] be a column of a unitary matrix?
[-i ]
Justify your answer
b. Check whether or not the following linear system is consistent:
3+t = x+y+z, 2+2t= x+2y-z, 4-t= x-y +4z
1
Expert's answer
2021-07-22T11:39:58-0400

Part a

Yes, [i] can be a column at a unitary matrix as if we choose "1 \times 1" matrix. I.e i=j, meaning that row = column.

So in this case we can say that [i] as a column of a unitary matrix.


Consqueantly

Yes, [-i] can be a column at a unitary matrix as if we choose "1 \times 1" matrix. I.e -i=-j, meaning that row = column.

So in this case we can say that [-i] as a column of a unitary matrix.


Part b

"x+y+z =3+t \\
 x+2y-z=2+2t\\
 x-y +4z=4-t\\"

It can be written as

"x+y+z -t=3 \\
 x+2y-z-2t=2\\
 x-y +4z+t=4\\"

Now we can convert it into matrix [A:b] form and do row reduction form

"=\begin{bmatrix}
 1 & 1 & 1 & -1| 3 \\
 1 & 2 & -1 & -2| 2\\
 1 & -1 & 4 & 1| 4
\end{bmatrix}\\"

"=\begin{bmatrix}
 1 & 1 & 1 & -1| 3 \\
 0 & 1 & -2 & -1| -1\\
 0 & -2 & 3 & 2| 1
\end{bmatrix}"


"=\begin{bmatrix}
 1 & 0 & 3 & 0| 4 \\
 0 & 1 & -2 & -1| -1\\
 0 & 0 & -1 & 0| -1
\end{bmatrix}"


"=\begin{bmatrix}
 1 & 0 & 0 & 0| 1 \\
 0 & 1 & 0 & -1| 1\\
 0 & 0 & 1 & 0| 1
\end{bmatrix}"

x= 1, y=1,z= 1

So the solution is consistent, the value of t is not fixed as t is a free variable. So it has infinitely many solutions.